OpenStudy (anonymous):

What is 5.6 x 10^-3yd/year in feet per day? answers: 6.813 x 10^-7 6.132 x 10^-6 5 x 10^-6 4.6 x 10^-5

OpenStudy (campbell_st):

so convert to feet by multiplying by 3 so 3 * 5.6 x 10^-3 feet/year 1 year is 365 days so it can be written as \[3 \times 5.6 \times 10^{-3} feet/365 days\] divide by 365 to get feet per day \[\frac{3 \times 5.6 \times 10^{-3}}{365}~~ feet/day\] just calculate it out

OpenStudy (anonymous):

Thanks